# Unveiling Over-Memorization in Finetuning LLMs for Reasoning Tasks

### v1

@@ 7.1 Recipes of Checkpoint Selection
To address the above issues, we offer the following suggestions: (1) When finetuning LLMs, it is advisable to use a small number of epochs, typically between 1 and 4. This reduces the likelihood of over-memorization, allowing the last checkpoint to be used as the final model. (2) For model selection, it is recommended to use a combination of validation accuracy and perplexity, choosing models with the highest validation accuracy within a certain range of validation perplexity. Alternatively, checkpoint averaging (Schmidt et al., 2023; Herold et al., 2024) can also be used to obtain the final model, integrating the advantages of multiple checkpoints.
7.2 Recipes of Learning Rate Configuration
Our study finds that all finetuning methods are sensitive to the learning rate. This aligns with previous research (Lialin et al., 2024; Biderman et al., 2024), underscoring the importance of setting an appropriate learning rate when finetuning models. Specifically, we have the following recommendations for learning rate configuration: (1) Large learning rates increase the risk of over-memorization over the same training period, and an excessively large learning rate can alter the learning patterns from the very beginning of training, severely impacting model performance (Section 4.1). (2) The optimal learning rate correlates with the finetuning methods; full finetuning requires a smaller learning rate compared to PEFT methods like LoRA, according to the test performance in Section 4.3. (3) The learning rate is related to the model size and dataset size; the larger the model and dataset, the smaller the optimal learning rate (Section 6.2). Therefore, careful tuning of the learning rate is crucial to achieve optimal and robust performance. We leave the exploration of more specific design rules for its configuration to future work.
@@ 8 Conclusion
In this paper, we introduce and systematically analyze the phenomenon of over-memorization in LLM finetuning, where test accuracy remains stable despite a rapid increase in test perplexity. We study the factors such as learning rates, training epochs, and finetuning methods that influence over-memorization. We analyze its impact on LLM model behavior, including negative effects on model robustness, OOD performance, calibration, generation diversity, and privacy protection. Our scaling analyses show that over-memorization is common regardless of the finetuning dataset scale or the LLM model size. Our findings enhance practitioners’ understanding of LLM finetuning, providing insights into the selection of checkpoints and learning rates to avoid over-memorization. Additionally, our research highlights that overparameterized pretrained LLMs possess unique properties distinct from traditional machine learning models. We encourage future research to further empirically and theoretically investigate the training and generalization mechanisms of LLMs.
Appendix A Impact Statements
Limitation
While our work introduces the "over-memorization" phenomenon in LLM finetuning, we acknowledge that our work has a limitation in that our analysis predominantly focused on reasoning tasks. Although supplementary AlpacaEval experiments were included, the exploration of this phenomenon in open-ended generation tasks remains less comprehensive and warrants further study.
Ethical Considerations
@@ Appendix A Impact Statements
Future Societal Consequences
Our work introduces the ’over-memorization’ phenomenon in LLM finetuning and outlines our investigation into its characteristics, causes, and impacts on model behavior. This significantly aids a more comprehensive understanding of the LLM finetuning process and the resulting models’ true capabilities, which is vital for enhancing AI reliability and trustworthiness. Crucially, this understanding, by revealing how over-memorization can mask underlying model flaws despite good surface-level accuracy, informs more robust model development and evaluation practices. Such improvements directly contribute to the training of more trustworthy and stronger models in the wild.
@@ B.1 Dataset Setup
Unless otherwise specified, this paper utilizes LLaMA-3.1-8B as the base language model, with the first 100K samples from the MetaMathQA dataset serving as the training set. A subset of this data, denoted as MetaTrain, is first sampled to monitor training dynamics such as perplexity and accuracy. An additional set, MetaTest, is constructed from the remaining, unseen portion of MetaMathQA. Both MetaTrain and MetaTest are set to have the same sample size as the GSM8K test set. Since MetaTest originates from the same distribution as the training data, it is inherently more similar to the training set than GSM8K. Therefore, for clarity and ease of distinction, we designate MetaTest as the validation set and GSM8K as the test set by default. MetaMathQA is derived from GSM8K and MATH, making all three datasets share similar formats and reasoning styles. As a result, in Section 5.1, we treat MetaTest, GSM8K, and MATH as in-distribution (ID) test sets.
@@ B.3 Prompt Format
Appendix C Additional Behavioral Analyses of Over-memorized Models
The main paper (Section 5) discusses various behavioral aspects of over-memorized models. Due to space constraints in the main text, the detailed evaluations of model calibration (Appendix C.1) and membership inference attack performance (Appendix C.2) are presented in this section. These analyses provide further insights into the characteristics of over-memorized models.
C.1 Model Calibration
@@
C.2 Membership Inference Attack Performance
@@
C.3 More Diversity Metric
@@
Appendix D Supplementary Experimental Results
D.1 Effect of Training Dataset Size
The experiments results of effect of training dataset size is shown in Figure 8.
@@ D.1 Effect of Training Dataset Size
D.2 Additional Open-ended Task
@@
D.3 Additional Experiments for Different Finetuning Methods
@@
Regarding the model size scaling experiments with Qwen2.5 models, introduced in Section 6.2, the main paper included test set perplexity and accuracy for four learning rates, with figures typically visualizing results for one learning rate across different model sizes. To facilitate a clearer analysis of how each individual model’s performance varies across these different learning rates, this appendix presents the test set results in an alternative format in Figure 14. In these figures, each subfigure is dedicated to a single model size (e.g., Qwen2.5-0.5B) and displays its performance across the aforementioned learning rates.
D.4 Additional Experiments for Code Generation

### v2

@@ 6.1 Receipt for Checkpoint Selection
To address the above issues, we offer the following suggestions: (1) When finetuning LLMs, it is advisable to use a small number of epochs, typically between 1 and 4. This reduces the likelihood of over-memorization, allowing the last checkpoint to be used as the final model. (2) For model selection, it is recommended to use a combination of validation accuracy and perplexity, choosing models with the highest validation accuracy within a certain range of validation perplexity.
6.2 Mitigating Over-Memorization via Checkpoint Merging
Our first strategy mitigates over-memorization by merging checkpoints from different training stages. Earlier checkpoints (e.g., epoch 3) tend to preserve higher output diversity and OOD performance, whereas later checkpoints (e.g., epoch 10) achieve stronger ID accuracy but suffer from reduced diversity and OOD accuracy, as discussed in Section 5. By merging them, we aim to combine these complementary strengths, effectively ensembling the capabilities of different training stages (Dang et al., 2025; Wortsman et al., 2022).
Concretely, we follow the training and evaluation setup in Section 5, where we finetune LLaMA-3.1-8B with LoRA (learning rate $2\times 10^{-4}$) on MetaMath-100K and subsequently evaluate the model’s ID/OOD performance, as well as its generation diversity. Let $\theta_{\text{base}}$ denote the pre-trained parameters, $\theta_{3}$ the parameters after epoch 3, and $\theta_{10}$ those after epoch 10. We compute deltas $\Delta_{3}=\theta_{3}-\theta_{\text{base}}$ and $\Delta_{10}=\theta_{10}-\theta_{\text{base}}$, and define the merged model as:
$\theta_{\text{merge}}=\theta_{\text{base}}+\tfrac{1}{2}(\Delta_{3}+\Delta_{10}).$ (1)
As shown in Figure 6(a) and Table 5, checkpoint merging gets the best Pass@1 across all checkpoints and also has a great Pass@10 accuracy. Additionally, SFT Merge achieves a great average OOD accuracy, surpassing epoch 3 and epoch 10 by 1.2 and 3.3 points, respectively. These results indicate that merging effectively alleviates over-memorization without requiring additional training.
6.3 Mitigating Over-Memorization via Memorization-Aware Reweighting
(a) Pass@N comparison across methods.
(b) Accuracy and perplexity curves during training.
Figure 6: Evaluation of over-memorization mitigation methods. (a) Pass@N comparison across methods. (b) Accuracy (left) and perplexity (right) during 10 training epochs.
Table 5: Accuracy of methods on in-domain (ID) and out-of-domain (OOD) benchmarks. Avg. denotes the average over OOD test sets. Detailed descriptions of OOD benchmarks are in Section 5.1.
Methods ID Test Set OOD Test Set
GSM8K MATH SVAMP ASDiv MAWPS TabMWP Minerva MMLU_STEM Avg.
SFT Epoch 3 75.6 28.9 77.1 81.7 89.7 67.1 27.4 17.9 60.2
SFT Epoch 10 76.4 28.5 75.3 79.1 89.6 63.6 29.2 12.0 58.1
SFT Merge 78.5 31.4 77.8 81.5 90.4 68.9 31.6 17.9 61.4
MAR Epoch 3 76.7 29.2 74.4 81.2 90.1 68.9 29.2 17.1 60.2
MAR Epoch 10 76.4 29.6 74.5 80.8 90.2 68.7 29.6 21.4 60.9
MAR Merge 77.1 31.2 76.7 82.0 91.4 68.4 33.2 23.5 62.5
While merging provides a training-free solution, we further mitigate over-memorization by modifying the finetuning objective itself. Standard SFT assigns equal weight to all tokens, regardless of their learning difficulty. As a result, frequently memorized tokens dominate optimization, reinforcing rigid memorization rather than improving generalization.
We introduce Memorization-Aware Reweighting (MAR), which downweights tokens that the model already predicts with high confidence and upweights those with lower probabilities. Formally, for each token $y_{t}$, we compute its predicted probability $p_{\theta}(y_{t}\mid x,y_{<t})$ and reweight its contribution by $1-p_{\theta}(y_{t}\mid x,y_{<t})$:
$\mathcal{L}_{\text{MAR}}=\sum_{t=1}^{T}\big(1-p_{\theta}(y_{t}\mid x,y_{<t})\big)\cdot\ell(y_{t}),$ (2)
where $\ell(y_{t})$ is the standard cross-entropy loss. This adjustment encourages the model to focus on tokens that remain uncertain, reducing excessive memorization of already-learned patterns.
Figure 6(b) and Table 5 summarize the results. Compared with SFT, MAR maintains comparable performance across both ID and OOD benchmarks but achieves substantially lower perplexity throughout training, with less than half of SFT’s perplexity at epoch 10. On OOD benchmarks, MAR does not suffer from performance degradation due to additional training; in fact, its accuracy at epoch 10 even surpasses that of MAR at epoch 3 and SFT at the same epoch. Furthermore, unlike SFT, MAR’s Pass@N accuracy at epoch 10 exceeds that at epoch 3, further suggesting that additional training does not lead to over-memorization in MAR. In addition, merging MAR checkpoints from different training stages (Epoch 3 and Epoch 10) further improves performance, yielding the highest Pass@10 accuracy and strongest OOD results among all evaluated methods. These results demonstrate that our proposed methods can effectively mitigate over-memorization, which achieves better ID accuracy compared to SFT, while preserving strong output diversity and OOD performance (see MAR Epoch 10 and MAR Merge).
@@ 7 Conclusion
In this paper, we introduce and systematically analyze the phenomenon of over-memorization in LLM finetuning, where test accuracy remains stable despite a rapid increase in test perplexity. We study the factors such as learning rates, training epochs, and fine-tuning methods that influence over-memorization. We analyze its impact on LLM model behavior, including negative effects on model robustness, OOD performance, calibration, generation diversity, and privacy protection. Our scaling analyses show that over-memorization is common regardless of the finetuning dataset scale or the LLM model size. Our findings enhance practitioners’ understanding of LLM finetuning, providing insights into checkpoint selection and introducing methods to mitigate over-memorization. Additionally, our research highlights that overparameterized pretrained LLMs possess unique properties distinct from traditional machine learning models. We encourage future research to further empirically and theoretically investigate the training and generalization mechanisms of LLMs.
Ethics statement
@@
Reproducibility statement
This work primarily presents an analysis of the "over-memorization" phenomenon in LLM fine-tuning. Full details of the training setup, including dataset configuration, hyperparameters for each fine-tuning method, and prompt designs, can be found in Appendix B of the paper.
Appendix A LLM Usage
In the preparation of this paper, we only used large language models (LLMs) as an assistive tool for grammar correction and text polishing.
@@ B.1 Dataset Setup
Unless otherwise specified, this paper utilizes LLaMA-3.1-8B as the base language model, with the first 100K samples from the MetaMathQA dataset serving as the training set. A subset of this data, denoted as MetaTrain, is first sampled to monitor training dynamics such as perplexity and accuracy. An additional set, MetaTest, is constructed from the remaining, unseen portion of MetaMathQA. Both MetaTrain and MetaTest are set to have the same sample size as the GSM8K test set. Since MetaTest originates from the same distribution as the training data, it is inherently more similar to the training set than GSM8K. Therefore, for clarity and ease of distinction, we designate MetaTest as the validation set and GSM8K as the test set by default. MetaMathQA is derived from GSM8K and MATH, making all three datasets share similar formats and reasoning styles. As a result, in Section 5.1, we treat GSM8K and MATH as in-distribution (ID) test sets.
@@ B.3 Prompt Format
Appendix C Mechanistic Explanation of Over-Memorization
The paradoxical coexistence of high accuracy and high perplexity in over-memorized LLMs can be explained by the cross-entropy (CE) loss used in supervised finetuning. Given input $x$ and target sequence $y_{1:T}$,
$\mathcal{L}_{\mathrm{CE}}(\theta)=-\sum_{t=1}^{T}\log p_{\theta}(y_{t}\mid x,y_{<t}),$ (3)
with gradient
$\frac{\partial\mathcal{L}_{\mathrm{CE}}}{\partial z_{t}(j)}=p_{t}(j)-\mathbf{1}\{j=y_{t}\}.$ (4)
Here, the annotated token $y_{t}$ is consistently reinforced, while all alternatives $a\neq y_{t}$ are uniformly suppressed. Because $p_{t}(a)\propto e^{z_{t}(a)}$, repeated updates drive an exponential decay in their probabilities. The dominant reasoning path is preserved, but alternative paths become increasingly implausible.
This mechanism explains why the set of high-perplexity tokens remains stable between a normally finetuned model and an over-memorized model, yet their perplexity magnitudes escalate dramatically. Accuracy remains unaffected because the dominant path is still valid, while perplexity rises due to systematic discouragement of all other continuations.
Appendix D Additional Behavioral Analyses of Over-memorized Models
The main paper (Section 5) discusses various behavioral aspects of over-memorized models. Due to space constraints in the main text, the detailed evaluations of model calibration (Appendix D.1) and membership inference attack performance (Appendix D.2) are presented in this section. These analyses provide further insights into the characteristics of over-memorized models.
D.1 Model Calibration
@@
D.2 Membership Inference Attack Performance
@@
D.3 More Diversity Metric
@@
Appendix E Supplementary Experimental Results
E.1 Scaling Analysis
Effect of Model Size
(a) (a) 2e-6
(b) (b) 2e-5
(c) (c) 2e-4
(d) (d) 5e-4
@@ E.1 Scaling Analysis
We further investigate the effect of model scale on the over-memorization phenomenon using Qwen2.5 models (Team, 2024b) of varying sizes (0.5B, 1.5B, and 3B), with all finetuning performed via LoRA, consistent with our main experimental setup. Figure 8 reveals several key trends: (1) Larger models generally yield better test performance. (2) Over-memorization is more pronounced in larger models, as indicated by substantially higher test perplexity in the overtraining regime (Figures 8(c)–(d)). (3) Larger models exhibit stronger initial generalization, achieving lower test perplexity early in training. However, their greater capacity for memorization (Tirumala et al., 2022) leads to a faster increase in test perplexity over time, making them more susceptible to over-memorization (Figures 8(c)–(d)). (4) Larger models favor smaller optimal learning rates, as shown in Figure 15 (Appendix E). This may be due to the aggregate effect of conservative updates across a large number of parameters, enabling effective optimization even with small per-parameter steps.
Effect of Data Size
The results in Figure 9 reveal several key insights: (1) The over-memorization pattern—characterized by rising test perplexity despite stable accuracy—persists across all dataset sizes. (2) In line with prior work (Zhang et al., 2024a), larger datasets generally yield better test performance. (3) Most notably, we observe a systematic interaction between dataset size and optimal learning rate: larger datasets favor smaller learning rates. For instance, the optimal learning rate shifts from 2e-4 on the 5K subset to 2e-5 on the 100K subset, with the 30K subset performing well under both. This behavior likely stems from the increased number of optimization steps induced by larger datasets under fixed training epochs and batch size, making smaller learning rates sufficient to traverse the solution space.
E.2 Additional Open-ended Task
@@
E.3 Additional Experiments for Different Finetuning Methods
@@
Regarding the model size scaling experiments with Qwen2.5 models, introduced in Section E.1, the main paper included test set perplexity and accuracy for four learning rates, with figures typically visualizing results for one learning rate across different model sizes. To facilitate a clearer analysis of how each individual model’s performance varies across these different learning rates, this appendix presents the test set results in an alternative format in Figure 15. In these figures, each subfigure is dedicated to a single model size (e.g., Qwen2.5-0.5B) and displays its performance across the aforementioned learning rates.
E.4 Additional Experiments for Code Generation